How do I begin to solve these equations? 1) 1/2x+5/8=3/4x 2) 2x/4+x/5=7/10 3) -4(n+5)+3(2n-4)=5(n-4)
Found 2 solutions by unlockmath, mtdddog75:
Answer by unlockmath(1688)

Hello,
I'll make a deal with you. I'll do the first one and see if you can do the others. OK?
1) 1/2x+5/8=3/4x This one is best done by getting rid of the fractions first. So multiply by 8. I'm assuming the equation is actually this: 1) 1/2(x)+5/8=3/4(x)
4x+5=6x Now subtract 4x from both sides to get:
5=2x Divide each side by 2 to result in:
x=5/2 There we go.
The whole objective in algebra is to get the x on one side by itself.
RJ
Check out a new book I wrote at:
www.math-unlock.com

Answer by mtdddog75(2)

To begin solving these problems you need to bring the variable to the left side of the equation.
1/2x + 5/8= 3/4x
-5/8 -5/8 subtract 5/8 from each side
1/2x = 3/4x - 5/8
-3/4x -3/4x next subtract 3/4x from both sides
-1/4x = 5/8 next multiply each side by -4
x=5/2

checking put 5/2 into equation and you will get 15/8 = 15/8 so this is true

Mike